A Level Mathematics 9709 Paper 61, May/June 2025

Question paper and mark scheme50 marks7 questions

Solve it on ExamPal

Practise

Untimed. Write on the paper and check as you go.

Start

Exam

Timed, with a clock you set, like sitting the real paper.

Start

Opens full screen.

Opening the question paper

Read as textThe questions and the mark scheme as plain text. Easier to read on a phone.

Question paper

9709/61 · May/June 2025

Download PDF
1

It is known that 1% of houses in a certain area have a wind turbine. A random sample of 400 houses in this area is chosen for a survey on domestic heating. The number of houses in the sample that have a wind turbine is denoted by X. Use a suitable approximating distribution to find P(X⩽3).

[3]
2

The random variable X has the distribution B(8,34). A random sample of 100 values of X is chosen, and the sample mean, Xˉ, is found.

a

Find P(Xˉ>6.2). You are not expected to use a continuity correction.

[6]
b

State why the Central Limit Theorem was needed in the calculation in part (a).

[1]
3

The time, T minutes, for a certain daily bus journey is normally distributed. The bus company claims that the mean of T is 45. A passenger believes that the mean of T is actually greater than 45. She notes the times taken for this journey on a random sample of 60 days. The results are summarised below. n=60Σt=2750Σt2=127 000

a

Calculate unbiased estimates of the population mean and variance.

[3]
b

Test the passenger's belief at the 5% significance level.

[5]
4

At an entertainment centre, the cost for using a particular video game is $0.40 per minute. The number of minutes for which people use the video game has mean 15 and variance 9.

a

Find the mean and variance of the amount people pay for using the video game.

[3]
b

Each day, 35 people independently use the video game. Find the mean and variance of the total amount paid by 35 people.

[3]
5
a(i)

The random variables W and X have the independent distributions Po(1.2) and Po(2.3) respectively. Find P(3⩽W+X⩽5).

[2]
a(ii)

The random variable S is the sum of 100 independent values of W and 200 independent values of X. Use a suitable approximation to find P(S>600).

[6]
b

The random variable Y has the distribution Po(λ), where λ>0. It is given that 52P(Y=3)+P(Y=4)=P(Y=5). Find the value of λ.

[3]
6

A manufacturer of cell phones claims that 25% of students own a Pumpkin phone. Jeyeraj thinks that the proportion of students at his large college who own a Pumpkin phone is less than 25%. He plans to test the manufacturer's claim. He chooses a random sample of 30 students at his college. If the number of students who own a Pumpkin phone is less than 5, Jeyeraj will reject the manufacturer's claim.

a

State suitable hypotheses for the test.

[1]
b

Given that the true proportion of students at the college who own a Pumpkin phone is 10%, use a binomial distribution to find the probability of a Type II error.

[3]
c

At Florence's college, in a random sample of 40 students, it was found that 5 own a Pumpkin phone. Calculate an approximate 95% confidence interval for the proportion of students at Florence's college who own a Pumpkin phone.

[3]
7

X is a random variable with probability density function given by f(x)={(1+cos⁡πx)0⩽x⩽1,0otherwise.

a

Show that P(X<12)=12+1π.

[3]
b

Show that E(X)=12−2π2.

[5]

Mark scheme

9709/61 · May/June 2025

Download PDF

1 · 3 marks

  • B1

    (λ) = 4

  • M1

    e−4(1+4+422!+433!)=e−4(1+4+8+10.6666)=0.018316+0.073263+0.146525+0.19537

  • A1

    = 0.433

2a · 6 marks

  • B1

    N(6, ....)

  • M1

    σ2=32=1.5 (or σ=1.5 or 62 or 1.22 (3sf))

  • M1

    6.2−61.5100

  • A1

    = 1.633

  • M1

    P(X‾>6.2)=1−Φ(‘1.633’)

  • A1

    = 0.0512 or 0.0513 (3sf)

2b · 1 mark

  • B1

    Population (X) is not normally distributed / Population (X) is Binomial

3a · 3 marks

  • B1

    Est(μ)=275060 or 2756 or 45.8 (3 sf)

  • M1

    Est(σ2)=6059(12700060−(2756)2)

  • A1

    = 16.2 (3 sf) or 2875177

3b · 5 marks

  • B1

    H0:Population mean time=45H1:Population mean time>45

  • M1

    ‘2756’−45‘16.24294’60

  • A1

    = 1.602 to 1.595

  • M1

    ‘1.602’ < 1.645

  • A1

    [Accept H0] There is insufficient evidence [at 5% level] to reject the company’s claim OR There is insufficient evidence to accept the passenger’s belief OR There is insufficient evidence that the mean time is more than 45 minutes

4a · 3 marks

  • B1

    Mean = 6 [$]

  • M1

    Variance = 0.402×9

  • A1

    = 1.44 [$2]

4b · 3 marks

  • B1

    Mean = 35ב6’=210 [$]

  • M1

    Variance = 35ב1.44’

  • A1

    = 50.4 [$2]

5a(i) · 2 marks

  • M1

    e−3.5(3.533!+3.544!+3.555!)=e−3.5(7.1458+6.2526+4.37682)=0.21579+0.18881+0.13217

  • A1

    = 0.537

5a(ii) · 6 marks

  • B1

    Use of N(1.2×100+2.3×200,… )=N(580,… )

  • M1

    Var(S)=1.2×100+2.3×200=580

  • M1

    600.5−‘580’‘580’

  • A1

    = 0.851

  • M1

    1−Φ(‘0.581’)

  • A1

    = 0.197 (3sf)

5b · 3 marks

  • B1

    52×λ33!+λ44!=λ55! Or (52)e−λ(λ33!)+e−λ(λ44!)=e−λ(λ55!)

  • M1

    λ2−5λ−50=0

  • A1

    (λ−10)(λ+5)=0 or λ=5±25+4×502 λ=10

6a · 1 mark

  • B1

    H0:Proportion (at college) owning Pumpkin phone=0.25 H1:Proportion (at college) owning Pumpkin phone<0.25

6b · 3 marks

  • M1

    B(30,0.1) and P(X⩾5) attempted

  • M1

    1−(0.930+30×0.929×0.1+30C2×0.928×0.12+30C3×0.927×0.13+30C4×0.926×0.14)=1−(0.042391+0.141304+0.22766+0.236088+0.177066)

  • A1

    = 0.175 (3 sf) accept 0.176

6c · 3 marks

  • M1

    18±z18×7840

  • B1

    z=1.96

  • A1

    0.0225 to 0.227 (3 sf)

7a · 3 marks

  • M1

    ∫012(1+cos⁡πx) dx

  • A1

    =[x+sin⁡πxπ]012

  • A1

    =12+sin⁡π2π=12+1π

7b · 5 marks

  • M1

    ∫01(x+xcos⁡πx) dx

  • M1

    =[x22+xsin⁡πxπ−∫sin⁡πxπ dx]01

  • A1

    =[x22+xsin⁡πxπ+cos⁡πxπ2]01

  • 1 mark

    =12+sin⁡ππ+cos⁡ππ2−cos⁡0π2

  • A1

    =12−2π2

Questions

  1. 1
    Question 13 marks
    Mark scheme ›
  2. 2
    Question 27 marks
    Mark scheme ›
  3. 3
    Question 38 marks
    Mark scheme ›
  4. 4
    Question 46 marks
    Mark scheme ›
  5. 5
    Question 511 marks
    Mark scheme ›
  6. 6
    Question 67 marks
    Mark scheme ›
  7. 7
    Question 78 marks
    Mark scheme ›

More 9709 papers

Other papers from May/June 2025

Solve it on ExamPalStart